OpenStudy (automaticloveletter):

Two equations are given below: a − 3b = 4 a = b − 2 What is the solution to the set of equations in the form (a, b)? (−2, −2) (−3, −1) (−9, −7) (−5, −3)

OpenStudy (anonymous):

(b-2) - 3b = 4 b-2 - 3b = 4 -2b = 6 b = -3 a - 3(-3) = 4 a + 9 = 4 a = -5 Your answer is D ( -5, -3) Does it helps?

OpenStudy (automaticloveletter):

Thank you for showing your work.